OpenStudy (anonymous):

A New York Volleyball Association invited 64 teams to compete in a tournament. After each round, half the teams were eliminated. Whats the equation that represents the number of teams t, that remained in the tournament after r rounds?

OpenStudy (anonymous):

t= 64^(1/2r)

OpenStudy (anonymous):

err That should be not raised but multiplied

OpenStudy (anonymous):

no... that's still wrong. t= 64 * (1/2^r) is right

OpenStudy (anonymous):

yup, multiplied

OpenStudy (anonymous):

no, it's multiplied exponent doesn't make sense

OpenStudy (anonymous):

it makes complete sense, you need the exponent since each round you are multiplying by 1/2

OpenStudy (anonymous):

theres 4 answers sry bout that.... (1)t=64(r)^0.5 (3)t=64(1.5)^r (2)t=64(-0.5)^r (4)t=64(0.5)^r

OpenStudy (anonymous):

no wait you're right you need exponents it should be t=64(1/2)^r i take back what i said

OpenStudy (anonymous):

That's number 4 btw lilgumbunny

OpenStudy (anonymous):

Kk thx a bunch!